Find the volume V of the solid obtained by rotating the region bounded by the given curves about the specified line.

y = 4, sqrt49 − x2 , y = 0, x = 5, x = 6; about the x-axis.
Answer #1

25. y=4sqrt(49-x^2) x=6 20 15 10. 5. -5 0 5 10 y=0 X=5 5 《Given, x=6 y=4749-22 y=0, 2=5 TO colculate volume obtained by ratating the region bounded by these Cuzye when rotated about x